OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic charged $75 per hour, and the second mechanic charged $110 per hour. The mechanics worked for a combined total of 20 hours, and together they charged a total of $2025. How long did each mechanic work? Will medal

OpenStudy (jdoe0001):

so.. let's say hmmm the hours the $75 mechanic worked was \(\bf x\) and the hours the $110 mechanic worked was \(\bf y\) how many hours did they work together?

OpenStudy (anonymous):

20

OpenStudy (jdoe0001):

so... they worked together 20 hours we dunno what those hours are, but we can call them "x" and "y" so we could say that \(\bf x+y=20\) the mechanic that charges $75 per hour, after "x" hours, he ended up charging 75*x or 75x the mechanic that charges $110 per hour, after "y" hours, he ended up charging 110*y or 110y now, their total bill was 2025 so we could also say that \(\bf 75x+110y = 2025\) any ideas?

OpenStudy (anonymous):

im confused i have to give hours for mechanic one and mechanic 2

OpenStudy (anonymous):

i have 4 questions like this i really want to know how tp figure them out

OpenStudy (jdoe0001):

yes.... notice... x+y = 20 so... if you solve that for say "y"... how does it look like?

OpenStudy (jdoe0001):

the hours, notice above, for each is "x" and "y", thus

OpenStudy (texaschic101):

this is basically a substitution problem involving the 2 equations up above ^^. We have : x + y = 20 ----> x = -y + 20 75x + 110y = 2025 now sub -y + 20 in for x back into the second equation....you will then be solving for y, the hrs for the mechanic that charges 110 per hr.

OpenStudy (anonymous):

mechanic 1 5 hours mechanic 2 15 hours ??

OpenStudy (texaschic101):

lets check by subbing in your answers into the 2nd equation.. 75x + 110y = 2025 75(5) + 110(15) = 2025 375 + 1650 = 2025 2025 = 2025 (correct) yes...that is correct

OpenStudy (anonymous):

ty for explaining it

OpenStudy (texaschic101):

medal should go to jdoe....he/she helped you set it up...without that, you have nothing...lol

OpenStudy (anonymous):

you stayed to make sure i understood..

OpenStudy (texaschic101):

all is good...I gave jdoe a medal I am always happy to help when I can :)

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

:)